I have to solve
6x-2y=-22
4x+7y=2
The answers are (-3,2)

Answer by edjones(8007)

6x-2y=-22
4x+7y=2
.
-2y=-6x-22
y=3x+11
.
4x+7(3x+11)=2
4x+21x+77=2
25x=-75
x=-3
.
4*-3+7y=2
-12+7y=2
7y=14
y=2
.
Ed